# EmoVoice: LLM-based Emotional Text-To-Speech Model with Freestyle Text Prompting

### v3

@@ Abstract.
Human speech goes beyond the mere transfer of information; it is a profound exchange of emotions and a connection between individuals. While Text-to-Speech (TTS) models have made huge progress, they still face challenges in controlling the emotional expression in the generated speech. In this work, we propose EmoVoice, a novel emotion-controllable TTS model that exploits large language models (LLMs) to enable fine-grained freestyle natural language emotion control, and a phoneme boost variant design that makes the model output phoneme tokens and audio tokens in parallel to enhance content consistency, inspired by chain-of-thought (CoT) and chain-of-modality (CoM) techniques. Besides, we introduce EmoVoice-DB, a high-quality 40-hour English emotion dataset featuring expressive speech and fine-grained emotion labels with natural language descriptions. EmoVoice achieves state-of-the-art performance on the English EmoVoice-DB test set using only synthetic training data, and on the Chinese Secap test set using our in-house data. We further investigate the reliability of existing emotion evaluation metrics and their alignment with human perceptual preferences, and explore using SOTA multimodal LLMs GPT-4o-audio and Gemini to assess emotional speech. Demo samples are available at https://yanghaha0908.github.io/EmoVoice/. Dataset, code, and checkpoints will be released.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: arXiv Preprint; April, 2025; arXiv.orgccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
@@ 1. Introduction
In this paper, we explore the task of fine-grained emotion-controllable TTS models with freestyle text prompting, and address limitations of existing works from the perspectives of model design, data, and evaluation metrics. At the model level, we employ an LLM-based TTS model to enhance comprehension of fine-grained emotion descriptions in natural language, harnessing LLM’s capabilities in textual semantic understanding and emotion analysis. Our model allows direct input of emotion descriptions into the LLM, using the LLM to comprehend instructions without dedicated prompt encoders required by previous models (Guo et al., 2023b). Besides, inspired by the CoT (Wei et al., 2022) technique for LLMs and CoM proposed in SpeechGPT (Zhang et al., 2023), our model is designed to output phoneme tokens and audio tokens in parallel, forcing the model to predict pronunciation before generating the final audio tokens, hence boosting the performance. At the data level, we construct a high-quality synthetic emotion dataset by distilling the current best-performing TTS model, GPT-4o-audio [^1] https://platform.openai.com/docs/models/gpt-4o-audio-preview, which excels in emotional expressiveness and instruction-following capabilities. At the evaluation metric level, we delve into the reliability of existing emotion metrics and their alignment with human preferences, and pioneer the effort to evaluate emotional speech with SOTA multimodal LLMs GPT-4o-audio1 and Gemini [^2] https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash with comprehensive capabilities.
@@ 2.1. Emotional TTS
Coarse-grained emotion category. Earlier works mainly use category labels for emotion control, focusing on mainstream categories such as happy, sad, angry, etc. Some emotional TTS systems adopt diffusion models (Guo et al., 2023a; Kang et al., 2023; Tang et al., 2023a, 2024). Among them, EMODIFF (Guo et al., 2023a) allows manipulation of emotion intensity through soft-label. EmoMix (Tang et al., 2023a) combines a diffusion probabilistic model with a pre-trained speech emotion recognition model, enabling generation of emotional speech with specified intensity or blended emotions. (Tang et al., 2023b; Wang et al., 2023b; Cho et al., 2024b; Diatlova and Shutov, 2023) build on the non-autoregressive transformer-based FastSpeech2 model (Ren et al., 2021). For example, EmoSpeech (Diatlova and Shutov, 2023) introduces a conditioning mechanism to address uneven emotion distribution in text, enabling emotion intensity variation across phonemes for improved intonation perception. There are also explorations beyond model architecture. For example, Emo-DPO (Gao et al., [n. d.]) leverages an LLM-TTS architecture and direct preference optimization to capture subtle emotional nuances. EmoPro (Wang et al., 2024b) proposes a two-stage prompt selection strategy for selecting high-quality prompt speech.
@@
Fine-grained natural language description. Using natural language descriptions as emotion labels, such as Expressing supportive joy and pride in someone’s accomplishment, is superior to simple categorical tags by capturing complex and nuanced emotional states, allowing for richer emotional expression in speech synthesis. However, no existing TTS models are specifically designed to control emotions through natural language prompts. Most instead focus on general style control, adjusting attributes including gender, pitch, speed, volume, and emotion. For example, promptTTS (Guo et al., 2023b) conditions FastSpeech2 on style representations extracted from prompts with a style encoder. It trains on a custom dataset with five style factors—gender, pitch, speed, volume, and emotion. InstructTTS (Yang et al., 2024a) trains a discrete diffusion model to generate discrete acoustic features, and builds a Chinese corpus focusing on gender, pitch, speech, and volume. PromptTTS2(Leng et al., 2023) is a diffusion-based variation network modeling voice variability beyond prompts to address the one-to-many problem. Salle (Ji et al., 2024a) integrates style prompts into the Valle(Wang et al., 2023a) model. ControlSpeech (Ji et al., 2024b) is capable of simultaneously cloning a speaker’s voice based on an audio prompt and controlling speaking style with a textual style prompt. PromptStyle (Liu et al., 2023) achieves controllable style transfer through natural language prompts, while  (Bott et al., 2024) implicitly utilizes emotional semantics in an emotionally rich text as a prompt to control speech.
@@
The most relevant works to ours are CosyVoice (Du et al., 2024a) and CosyVoice2 (Du et al., 2024b). CosyVoice-Instruct, applies instruction fine-tuning on CosyVoice to control speaker identity, speaking style, and fine-grained paralinguistic features, based on hundreds of hours of internal data. Cosyvoice2 integrates 1500 hours of instructed training data into the base training set, seamlessly integrating the instruction and zero-shot capabilities into a single model. Our EmoVoice differs from CosyVoice2 in three key aspects. We employ group token modeling, predicting three audio tokens simultaneously at each decoding step, which significantly accelerates training and reduces modeling difficulty. Our EmoVoice-PP outputs phoneme tokens in parallel to guide audio token generation, which improves content consistency. Moreover, our models are trained on emotion-specific data of higher quality. Besides, OpenAI’s GPT-4o-audio remains the most advanced TTS model for natural language instruction control. The latest released GPT-4o-mini-TTS [^3] https://platform.openai.com/docs/guides/text-to-speechalso supports prompting to control various speech characteristics.
@@ 2.2. LLM-based TTS
LLMs have made significant advancements in natural language processing and are expanding their capabilities to speech and vision modalities. While most previous work concentrates on prompting LLMs with audio comprehension abilities, some studies also explore adapting LLMs for the TTS task. (Hao et al., 2025) presents a comprehensive empirical study on enhancing LLMs with speech generation capabilities by integrating pre-trained Llama/OPT (Touvron et al., 2023; Zhang et al., 2022) with VALL-E, comparing three integration methods. TTS-Llama (Shen et al., 2025) fine-tunes the Llama3-8B-Instruct with LoRA (Hu et al., 2021), being the first TTS system to achieve SOTA performance through only PEFT fine-tuning of a text-based LLM. Spark-TTS (Wang et al., 2025) is an innovative system powered by BiCodec, a single-stream speech codec that disentangles semantic and speaker information from speech, utilizing Qwen2.5-0.5B (Yang et al., 2024c) as backbone, integrating TTS capabilities within the text LLM paradigm. Llasa (Imran et al., 2024) is initialized from the Llama by expanding vocabulary to incorporate speech tokens, trained using the next-token prediction paradigm. It explores the scaling of train-time and inference-time compute for TTS task, and proposes a single-layer codec to better align with standard LLMs. CosyVoice2 upgrades CosyVoice by utilizing the Qwen2.5-0.5B to autoregressively generate speech tokens from text prompts, eliminating the need for a separate text encoder as the LLM is powerful enough to align the text and speech tokens. Our EmoVoice also adopts Qwen2.5-0.5B as the backbone. Besides, the Omni series models primarily employ LLM as the backbone to enable end-to-end speech conversation tasks. Mini-Omni (Xie and Wu, 2024a), Mini-Omni2 (Xie and Wu, 2024b), and SLAM-Omni (Chen et al., 2024) utilize the Qwen2-0.5B as their foundational model, while LLama-Omni (Fang et al., 2024) adopts Llama-3.1-8B-Instruct, and Freeze-Omni (Wang et al., 2024a) employs Qwen2-7B-Instruct3.
@@ 3.2. Training Pipeline
Figure 1. Overview of EmoVoice. LLM receives emotion description and text, autoregressively generating audio token group and phoneme token(optional) in parallel. Audio tokens are converted to waveforms by Vocoder. Input at each decoding step is average embedding of audio token group and phoneme token(optional).
@@ 4.1. Overview of EmoVoice-DB
We introduce EmoVoice-DB, an unprecedented, high-quality English emotional speech dataset featuring fine-grained emotion labels expressed through natural language descriptions. This dataset contains over 20,000 emotionally expressive speech samples, each annotated with detailed and precise emotional descriptions, totaling approximately 40 hours of audio. EmoVoice-DB is built using synthetic data generated by the powerful GPT-4o [^4] https://platform.openai.com/docs/models/gpt-4o and GPT-4o-audio models.
@@ 4.2. Dataset Construction Process
Step 1: Generating text and emotional descriptions. In the first step, pairs of texts and corresponding emotional descriptions are generated using the GPT-4o model. For each emotional category, the model is prompted to produce text that aligns with the specified emotion and a detailed emotional description that captures the intended sentiment. The text generation process is guided by a strict set of constraints to ensure quality and consistency. Each utterance is required to be between 15 and 25 words in length, corresponding to 5–10 seconds of speech. The text should contain rich sensory and emotional details while avoiding clichéd or overused patterns, such as “I can’t believe…”. To enrich the diversity and depth of the textual content, the texts are evenly distributed across three distinct categories: vivid descriptive sentences in the form of novel prose, e.g., “Wind WHISPERED through the parched cornstalks, its voice fraying like worn silk.”, emotionally charged dialog excerpts representing natural spoken lines, e.g., “I’ve ASKED you three times! Why is the door still locked?”, observational phrases offering subtle situational commentary, e.g., “Rain taps the window like it’s BRUISING the glass—rhythmic, insistent, all night.” To guide prosodic emphasis in the synthesized speech, up to two words per sentence could be capitalized to indicate stress or emphasis. We also impose specific constraints on the emotional description. Each description is formatted as a present participle verb phrase, e.g., “Conveying a contagious, joyful atmosphere.”, and should focus on vocal affect, excluding contextual references or specific events. In addition, descriptions are restricted to a single sentence, avoiding overly brief or single-word responses that lack sufficient details.
@@ 5. Experimental Setup
For the initial pretraining stage, we employ the VoiceAssistant (Xie and Wu, 2024a) dataset for the English TTS model and the Belle (Ji et al., 2023) dataset for the Chinese TTS model. VoiceAssistant and Belle are initially text dialogue datasets in English and Chinese, respectively. SLAM-Omni utilizes them to construct synthesized speech dialogue datasets. It employs the text-to-token LLM from CosyVoice to generate semantic tokens, with enriched vocal timbre. For training, we use only the response portions of the dialogues. The response portion of the VoiceAssistant dataset contains 3,234 hours of speech, while that of the Belle dataset consists of 6,418 hours of speech. We directly use the text and semantic token pairs as the training data.
@@
For the second fine-tuning stage, we employ two datasets for the English model: a synthetic emotional speech dataset by GPT-4o-audio with categorical labels, “LAION’s Got Talent,” and our proposed EmoVoice-DB, which features freestyle natural language descriptions of emotions. The “LAION’s Got Talent” dataset [^5] https://huggingface.co/datasets/laion/laions_got_talent consists of voice acting samples that span a broad range of emotions, vocal bursts, topics, and content. The emotional labels combine multiple synonyms to describe a specific emotion, such as ”intense anger, rage, fury, hatred, and annoyance, speaking without any accent.” These descriptions are fixed and finite, resembling emotion classification tags more than natural language descriptions. We filter the dataset to include only English sentences spoken without any accent, resulting in 44.4k samples, amounting to about 200 hours of speech. This filtered data is entirely used as the training set. Additionally, we randomly sample 100 utterances per emotion category from the EmoVoice-DB dataset for the test set and 50 utterances for the validation set, with the remaining audio serving as the training set. We utilize an internal industrial training dataset for the Chinese model and adopt the Secap (Xu et al., 2024) test set for evaluation, which contains a total of 600 audio samples. For all emotional fine-tuning data, we extract 50 Hz CosyVoice semantic tokens from each utterance. The resulting text-token pairs are directly used as input training data.
@@
Training and Inference Details. In the training process, EmoVoice (0.5B) is initialized with the pre-trained Qwen2.5-0.5B model and train all parameters for the speech synthesis tasks. The complete architecture, comprising the LLM and grouped decoding linear layer, contains 550M trainable parameters. EmoVoice(1.5B) utilizes Qwen2.5-1.5B of 28 Transformer layers with 1.54B parameters. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with $\beta=(0.9,0.999)$ and zero weight decay. The semantic token group size is set to 3, following SLAM-Omni, which enables 2.64x faster training than standard single-layer audio token output, with lower WER and higher content accuracy. During the standard TTS pre-training phase, the learning rate increases linearly from 0 to a peak rate of 1e-4 over the first 1,000 steps, and then decays linearly to zero during the remaining training time. In the emotional TTS fine-tuning stage, the learning rate scheduler follows the same pattern, but with the peak rate set to 1e-5. We conduct our experiments on 4 80GB A800 GPUs and set the batch size to 6.
In the inference process, we use greedy search decoding algorithm with a repetition penalty of 1.2. We employ flow matching model and HiFi-GAN vocoder of CosyVoice to convert semantic tokens into audio waveforms. We pick a different neutral-emotion speech sample of the same speaker with the ground-truth speech, as the prompt speech for the flow matching module to control the timbre of the synthesized audio.
@@
Evaluation Metrics. We evaluate the content consistency of synthesized speech using the Word Error Rate (WER), with transcription results obtained from the Whisper large-v3 (Radford et al., 2023) model. To assess emotional expressiveness, we adopt two metrics: Emotion Similarity and Recall Rate. Emotion Similarity is computed by extracting emotion embeddings from the synthesized speech using emotion2vec, and calculating the cosine similarity between these embeddings and those from the ground-truth speech. For Recall Rate, we also leverage the emotion2vec model to perform emotion classification on the generated speech. Due to insufficient training data and consequently low recognition accuracy of emotion2vec for the categories disgusted, fearful, and surprised, these emotions are excluded from the evaluation. Recall is computed as the proportion of correctly classified audio samples within each emotion category, and the final Recall Rate is the average value across all selected categories. In addition, we evaluate the naturalness and perceptual audio quality of the synthetic speech using the UTMOS score (Saeki et al., 2022).

### v4

@@ Abstract.
Human speech goes beyond the mere transfer of information; it is a profound exchange of emotions and a connection between individuals. While Text-to-Speech (TTS) models have made huge progress, they still face challenges in controlling the emotional expression in the generated speech. In this work, we propose EmoVoice, a novel emotion-controllable TTS model that exploits large language models (LLMs) to enable fine-grained freestyle natural language emotion control, and a phoneme boost variant design that makes the model output phoneme tokens and audio tokens in parallel to enhance content consistency, inspired by chain-of-thought (CoT) and chain-of-modality (CoM) techniques. Besides, we introduce EmoVoice-DB, a high-quality 40-hour English emotion dataset featuring expressive speech and fine-grained emotion labels with natural language descriptions. EmoVoice achieves state-of-the-art performance on the English EmoVoice-DB test set using only synthetic training data, and on the Chinese Secap test set using our in-house data. We further investigate the reliability of existing emotion evaluation metrics and their alignment with human perceptual preferences, and explore using SOTA multimodal LLMs GPT-4o-audio and Gemini to assess emotional speech. Dataset, code, checkpoints and demo samples are available at https://github.com/yanghaha0908/EmoVoice.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754829isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
@@ 1. Introduction
In this paper, we explore the task of fine-grained emotion controllable TTS models with freestyle text prompting, and address limitations of existing works from the perspectives of model design, data, and evaluation metrics. At the model level, we employ an LLM-based TTS model to enhance comprehension of fine-grained emotion descriptions in natural language, harnessing LLM’s capabilities in textual semantic understanding and emotion analysis. Our model allows direct input of emotion descriptions into the LLM, using the LLM to comprehend instructions without dedicated prompt encoders required by previous models (Guo et al., 2023b). Besides, inspired by the CoT (Wei et al., 2022) technique for LLMs and CoM proposed in SpeechGPT (Zhang et al., 2023), our model is designed to output phoneme tokens and audio tokens in parallel, forcing the model to predict pronunciation before generating the final audio tokens, hence boosting the performance. At the data level, we construct a high-quality synthetic emotion dataset by distilling the current best-performing TTS model, GPT-4o-audio [^1] https://platform.openai.com/docs/models/gpt-4o-audio-preview, which excels in emotional expressiveness and instruction-following capabilities. At the evaluation metric level, we delve into the reliability of existing emotion metrics and their alignment with human preferences, and pioneer the effort to evaluate emotional speech with SOTA multimodal LLMs GPT-4o-audio1 and Gemini [^2] https://ai.google.dev/gemini-api/docs/models##gemini-2.0-flash with comprehensive capabilities.
@@ 2.1. Emotional TTS
Coarse-grained emotion category. Earlier works mainly use category labels for emotion control, focusing on mainstream categories such as happy, sad, angry, etc. Some emotional TTS systems adopt diffusion models (Guo et al., 2023a; Kang et al., 2023; Tang et al., 2023a, 2024). Among them, EMODIFF (Guo et al., 2023a) allows manipulation of emotion intensity through soft-label. EmoMix (Tang et al., 2023a) combines a diffusion probabilistic model with a pre-trained speech emotion recognition model, enabling generation of emotional speech with specified intensity or blended emotions. (Tang et al., 2023b; Wang et al., 2023b; Cho et al., 2024b; Diatlova and Shutov, 2023) build on the non-autoregressive transformer-based FastSpeech2 model (Ren et al., 2021; Liu et al., 2023a). For example, EmoSpeech (Diatlova and Shutov, 2023) introduces a conditioning mechanism to address uneven emotion distribution in text, enabling emotion intensity variation across phonemes for improved intonation perception. There are also explorations beyond model architecture. For example, Emo-DPO (Gao et al., [n. d.]) leverages an LLM-TTS architecture and direct preference optimization to capture subtle emotional nuances. EmoPro (Wang et al., 2024b) proposes a two-stage prompt selection strategy for selecting high-quality prompt speech.
@@
Fine-grained natural language description. Using natural language descriptions as emotion labels, such as Expressing supportive joy and pride in someone’s accomplishment, is superior to simple categorical tags by capturing complex and nuanced emotional states, allowing for richer emotional expression in speech synthesis. However, no existing TTS models are specifically designed to control emotions through natural language prompts. Most instead focus on general style control, adjusting attributes including gender, pitch, speed, volume, and emotion. For example, promptTTS (Guo et al., 2023b) conditions FastSpeech2 on style representations extracted from prompts with a style encoder. It trains on a custom dataset with five style factors—gender, pitch, speed, volume, and emotion. InstructTTS (Yang et al., 2024a) trains a discrete diffusion model (Liu et al., 2025, 2024) to generate discrete acoustic features, and builds a Chinese corpus focusing on gender, pitch, speech, and volume. PromptTTS2(Leng et al., 2023) is a diffusion-based variation network modeling voice variability beyond prompts to address the one-to-many problem. Salle (Ji et al., 2024a) integrates style prompts into the Valle(Wang et al., 2023a) model. ControlSpeech (Ji et al., 2024b) is capable of simultaneously cloning a speaker’s voice based on an audio prompt and controlling speaking style with a textual style prompt.
PromptStyle (Liu et al., 2023b) achieves controllable style transfer through natural language prompts, while  (Bott et al., 2024) implicitly utilizes emotional semantics in an emotionally rich text as a prompt to control speech.
@@
The most relevant works to ours are CosyVoice (Du et al., 2024a, b). CosyVoice-Instruct, applies instruction fine-tuning on CosyVoice to control speaker identity, speaking style, and fine-grained paralinguistic features, based on hundreds of hours of internal data. CosyVoice2 integrates 1500 hours of instructed training data into the base training set, seamlessly integrating the instruction and zero-shot capabilities into a single model. Our EmoVoice differs from CosyVoice2 in three key aspects. We employ group token modeling, predicting three audio tokens simultaneously at each decoding step, which significantly accelerates training and reduces modeling difficulty. Our EmoVoice-PP outputs phoneme tokens in parallel to guide audio token generation, which improves content consistency. Our models are trained on emotion-specific data of higher quality. Besides, OpenAI’s GPT-4o-audio remains the most advanced TTS model for natural language instruction control. The latest GPT-4o-mini-TTS also supports prompting to control various speech characteristics.
@@ 2.2. LLM-based TTS
LLMs have made significant advancements in natural language processing and are expanding their capabilities to speech and vision modalities. While most previous work concentrates on prompting LLMs with audio comprehension abilities, some studies also explore adapting LLMs for the TTS task. (Hao et al., 2025) presents a comprehensive empirical study on enhancing LLMs with speech generation capabilities by integrating pre-trained Llama/OPT (Touvron et al., 2023; Zhang et al., 2022) with VALL-E, comparing three integration methods. TTS-Llama (Shen et al., 2025) fine-tunes the Llama3-8B-Instruct with LoRA (Hu et al., 2021), being the first TTS system to achieve SOTA performance through only PEFT fine-tuning of a text-based LLM. Spark-TTS (Wang et al., 2025) is an innovative system powered by a single-stream speech codec that disentangles semantic and speaker information from speech, utilizing Qwen2.5-0.5B (Yang et al., 2024c) as backbone, integrating TTS capabilities within the text LLM paradigm. Llasa (Imran et al., 2024) is initialized from the Llama by expanding vocabulary to incorporate speech tokens, trained using the next-token prediction paradigm. It explores the scaling of compute for TTS task, and proposes a single-layer codec to better align with standard LLMs. CosyVoice2 upgrades CosyVoice by utilizing the Qwen2.5-0.5B to autoregressively generate speech tokens from text prompts, eliminating the need for a separate text encoder as the LLM is powerful enough to align the text and speech tokens. Our EmoVoice also adopts Qwen2.5-0.5B as the backbone. Besides, the Omni series models primarily employ LLM as the backbone to enable end-to-end speech conversation tasks. Mini-Omnis (Xie and Wu, 2024a, b), and SLAM-Omni (Chen et al., 2024) utilize the Qwen2-0.5B as their foundational model, while LLama-Omni (Fang et al., 2024) adopts Llama-3.1-8B-Instruct, and Freeze-Omni (Wang et al., 2024a) employs Qwen2-7B-Instruct3.
@@ 3.2. Training Pipeline
Figure 1. Overview of EmoVoice. LLM receives emotion description and text, autoregressively generating audio token group and phoneme token(optional) in parallel.
@@ 4.1. Overview of EmoVoice-DB
We introduce EmoVoice-DB, an unprecedented, high-quality English emotional speech dataset featuring fine-grained emotion labels expressed through natural language descriptions. This dataset contains over 20,000 emotionally expressive speech samples, each annotated with detailed and precise emotional descriptions, totaling approximately 40 hours of audio. EmoVoice-DB is built using synthetic data generated by the powerful GPT-4o models.
@@ 4.2. Dataset Construction Process
Step 1: Generating text and emotional descriptions. In the first step, pairs of texts and corresponding emotional descriptions are generated using the GPT-4o model. For each emotional category, the model is prompted to produce text that aligns with the specified emotion and a detailed emotional description that captures the intended sentiment. The text generation process is guided by a strict set of constraints to ensure quality and consistency. Each utterance is required to be between 15 and 25 words in length, corresponding to 5–10 seconds of speech. The text should contain rich sensory and emotional details while avoiding clichéd or overused patterns. To enrich the diversity and depth of the textual content, the texts are evenly distributed across three distinct categories: vivid descriptive sentences in the form of novel prose, e.g., “Wind whispered through the parched cornstalks, its voice fraying like worn silk.”, emotionally charged dialog excerpts representing natural spoken lines, e.g., “I’ve asked you three times! Why is the door still locked?”, observational phrases offering subtle situational commentary, e.g., “Rain taps the window like it’s bruising the glass—rhythmic, insistent, all night.” To guide prosodic emphasis in the synthesized speech, up to two words per sentence could be capitalized to indicate stress. We also impose specific constraints on the emotional description. Each description is formatted as a present participle verb phrase, e.g., “Conveying a contagious, joyful atmosphere.”, and should focus on vocal affect, excluding contextual references or specific events. In addition, descriptions are restricted to a single sentence, avoiding overly brief or single-word responses that lack sufficient details.
@@ 5. Experimental Setup
For pretraining stage, we employ the VoiceAssistant (Xie and Wu, 2024a) dataset for the English model and the Belle (Ji et al., 2023) dataset for the Chinese model. VoiceAssistant and Belle are initially text dialogue datasets in English and Chinese, respectively. SLAM-Omni utilizes them to construct synthesized speech dialogue datasets. It employs the text-to-token LLM from CosyVoice to generate semantic tokens, with enriched vocal timbre. For training, we use only the response portions of the dialogues, containing 3,234 and 6,418 hours of speech from VoiceAssistant and Belle respectively.
@@
For fine-tuning stage, we employ two datasets for the English model: a synthetic emotional speech dataset by GPT-4o-audio with categorical labels, LAION’s Got Talent, and our proposed EmoVoice-DB, which features freestyle natural language descriptions of emotions. The LAION’s Got Talent dataset consists of voice acting samples that span a broad range of emotions, vocal bursts, topics, and content. The emotional labels combine multiple synonyms to describe a specific emotion, such as ”intense anger, rage, fury, hatred, and annoyance.” These descriptions are fixed and finite, resembling emotion classification tags more than natural language descriptions. We filter the dataset to include only English sentences with no accent, resulting in 44.4k samples, amounting to about 200 hours of speech. This filtered data is entirely used as the training set. Additionally, we randomly sample 100 utterances per emotion category from the EmoVoice-DB dataset for the test set and 50 utterances for the validation set, with the remaining audio serving as the training set. We utilize an internal industrial training dataset for the Chinese model and adopt the Secap (Xu et al., 2024) test set for evaluation, which contains a total of 600 audio samples. For all emotional fine-tuning data, we extract 50 Hz CosyVoice semantic tokens from each utterance. The resulting text-token pairs are directly used as input training data.
@@
Training and Inference Details. In the training process, EmoVoice (0.5B) is initialized with the pre-trained Qwen2.5-0.5B model and train all parameters. The complete architecture, comprising the LLM and grouped decoding linear layer, contains 550M trainable parameters. EmoVoice(1.5B) utilizes Qwen2.5-1.5B of 28 Transformer layers with 1.54B parameters. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with $\beta=(0.9,0.999)$ and zero weight decay. The semantic token group size is set to 3, which enables 2.64x faster training than standard single-layer audio token output, with lower WER and higher content accuracy. During the standard TTS pre-training phase, the learning rate increases linearly from 0 to a peak rate of 1e-4 over the first 1,000 steps, and then decays linearly to zero during the remaining training time. In the emotional TTS fine-tuning stage, the learning rate scheduler follows the same pattern, but with the peak rate set to 1e-5. We conduct our experiments on 4 80GB A800 GPUs and set the batch size to 6. In the inference process, we use greedy search decoding algorithm with a repetition penalty of 1.2. We employ flow matching model and HiFi-GAN vocoder of CosyVoice to convert semantic tokens into audio waveforms. We pick a different neutral-emotion speech sample of the same speaker with the ground-truth speech, as the prompt speech for the flow matching module to control the timbre of the synthesized audio.
@@
Evaluation Metrics. We evaluate the content consistency of synthesized speech using the Word Error Rate (WER), with transcription results obtained from the Whisper large-v3 (Radford et al., 2023) model. To assess emotional expressiveness, we adopt two metrics: Emotion Similarity and Recall Rate. Emotion Similarity is computed by extracting emotion embeddings from the synthesized speech using emotion2vec, and calculating the cosine similarity between these embeddings and those from the ground-truth speech. For Recall Rate, we also leverage emotion2vec to perform emotion classification on the generated speech. Due to insufficient training data and consequently low recognition accuracy of emotion2vec for the categories disgusted, fearful, and surprised, these emotions are excluded from the evaluation. Recall is computed as the proportion of correctly classified audio samples within each emotion category, and the final Recall Rate is the average value across all selected categories. In addition, we evaluate the naturalness and perceptual audio quality of the synthetic speech using the UTMOS score (Saeki et al., 2022).